OpenStudy (anonymous):

Integrate 2/(7y^2+4) dy

OpenStudy (anonymous):

please show me your work. I know is ARCTAN..but how do i get rid of the 4

OpenStudy (sirm3d):

\[\large \frac{ 2 }{ 7y^2+4 }=\frac{ 2 }{ 4 }\frac{ 1 }{ \left( \frac{ 7 }{ 4 } \right)y^2+1 }=\frac{ 1 }{ 2 }\frac{ 1 }{ \left( \frac{ \sqrt{7} }{ 2 }y \right) +1}\]

OpenStudy (sirm3d):

oopsie, it's \[\large \frac{ 1 }{ 2 }\frac{ 1 }{ \left( \frac{ \sqrt{7} }{ 2 }y \right)^2+1 }\]

OpenStudy (anonymous):

are you able to draw? because I can't seem to know all the symbols []/.etc